Does resonating molecule is always more stable than any non resonating molecule. Explain with example?

Grade:

3 Answers

vikas askiitian expert
13 years ago

deaR pranshu , the molecule where resonance  is possible is always more stable than a molecule in

which resonance is not possible ...

example - CH2=CH-CH=CH2(di-ene) is more stable than CH3CH=CH-CH3 due to possiblity of resonance...

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

The resonance energy of a compound is a measure of the extra stability of the conjugated system compared to the corresponding number of isolated double bonds.  This can be calculated from experimental measurements.

The resonance energy of benzene can be determined by comparing the enthalpies of hydrogenation of benzene and cyclohexene. If benzene did have the hypothetical 'cyclohexatriene' structure (the Kekulé structure), then it would be expected to have an enthalpy of hydrogenation that is three times that of cyclohxene.


The enthalpy of hydrogenation of cyclohxene is -121 kJ mol-1, providing a theoretical value of -363 kJ mol-1 for Kekulé benzene with its three separate carbon-carbon double bonds.

Gurwinder Kaur
13 years ago

Yes,bcoz rasonance effect  always dominates over any other effect.